# Optimization Proxies using Limited Labeled Data and Training Time – A Semi-Supervised Bayesian Neural Network Approach

### v1

@@ 5 Numerical Results and Discussion
Tables 1 and 2 present the comparative performance of different methods for solving the ACOPF problem on the ‘case57’ and ‘case118’ test cases, containing 57 and 118 nodes, respectively. For ’case57’, the Sandwich BNN SvP method achieves the best Max Eq. performance (0.027), outperforming all other methods, including the standard Sandwich BNN, without compromising other metrics. All the methods proposed in this paper outperform the best DNN results, typically achieved with the LD+MAE model (last row in the tables). Similar trends are observed for the ‘case118’ as well. [^6] See Appendix C for similar tabular results and discussion of trends on larger test instances.
It is important to contextualize the significance of these numerical improvements. In ACOPF problems, cost values are in USD, with the mean cost for ‘case118’ being \$97,000 or 9.7 in the per-unit system. Therefore, a 1% Gap corresponds to an expected difference of \$970 across the testing instances. A ‘Max Eq.’ value of 0.08 implies a maximum expected power imbalance of 8.0 Megawatts among all 118 nodes of ‘case118’. Thus, reducing the Max Eq. from 1.284 with the LD+MAE model to 0.089 with our Sandwich BNN SvP model represents a significant improvement.
@@
Figure 4: Comparison of voltage magnitude and voltage angle error bounds (in logarithmic scale) across bus indices for ‘case57’ (top row) and ‘case118’ (bottom row). The plot illustrates that PCBs using theoretical Bernstein bounds with 2 $\times$ MPV from hypothesis equation 5 are tightest among all PCBs. We consider $\delta=0.95$ and 1000 out-of-sample testing data points i.e. $M=1000$.
@@
Figure 2 illustrates the growth of various metrics with increasing system size while keeping training resources constant. The proposed BNN methods exhibit significantly lower scaling in the expected maximum equality gap. Although Sandwich BNN SvP shows slightly higher scaling inequality gaps, the relative improvement in Max Eq. far outweighs these minor drawbacks. Notably, for ‘case500’ (see Table 6 in the Appendix C), the expected maximum power imbalance is below 0.5% of the mean real power demand ($1.7\times 10^{4}$ MW) and power grids already are equipped with spinning reserves (see Ela et al. (2011)) that have reserve capacity to handle these imbalances. This is a significant improvement over the DNN models which have much higher ‘Mean Eq.’ values. Moreover, the ‘Max Ineq.’ growth could be easily suppressed by incorporating bound repair layers, as used in DNN models in ML4OPF AI4OPT (2023).
@@
Next, we present results for Proabilistic Confidence bounds, described in Section 4. Figure 3 shows that $2\cdot\text{MPV}$ consistently serves as an empirical upper bound for total variance in error, validating its robustness across models and system configurations [^7] Total variance in error is assumed to stabilizing with 1000 testing samples (see Figure 6 in Appendix C).. In Figure 4, the theoretical Bernstein bounds using $\mathbb{V}_{e}=2\cdot\text{MPV}$ provide tight, practical bounds, whereas Hoeffding’s bounds are overly conservative and not useful for grid operations. For example, in ‘case118’, the Bernstein bound ensures a probabilistic guarantee on voltage constraint satisfaction, such that a predicted voltage between 0.91–1.09 pu guarantees no violations within the ACOPF limits of 0.90–1.10 pu, i.e., the maximum value of the Bernstein bound on the error is 0.010 pu across all nodes. Compared to Hoeffding’s bound ($0.064$ pu) or the empirical Bernstein bound ($0.018$ pu), the Bernstein bound ($0.010$ pu) is far tighter and more practical, highlighting the benefits of BNNs for optimization proxies. The error bounds for the ‘case500’ is provided in Fig. 6 of the Appendix C. Finally, we note that both Hoeffding’s and empirical Bernstein bounds can also be obtained by testing DNN models across $M$ samples [^8] A form of MPV can be obtained via Ensemble DNNs Ganaie et al. (2022), however, it will lead to very high computational requirement compared to the BNN..
@@ 6 Conclusions
Acknowledgments
Use unnumbered third level headings for the acknowledgments. All acknowledgments, including those to funding agencies, go at the end of the paper.
@@ Appendix A ACOPF Problem: Modeling and Dataset
The alternating current optimal power flow (ACOPF) problem is essential for power grid operations and planning across various time scales. It determines generator set-points for real and reactive power that minimize generation costs while meeting power demand and satisfying physical and operational constraints on output variables. We follow the ACOPF model given in PowerModels Coffrin et al. (2018), and take the dataset from Torch_geometric (Lovett et al. (2024)).
@@ B.2.1 ML4OPF:
All other hyperparameters are set at default ML4OPF values AI4OPT (2023).
@@ B.3 Structure of posterior prediction matrix (PPM)
$\mathbf{Y}\equiv\begin{bmatrix}y_{11}&\cdots\cdots\cdots&y_{1H}\\ \vdots&\ddots\ddots\ddots&\vdots\\ \cdots&y_{\texttt{Variable,Sample}}&\cdots\\ \vdots&\ddots\ddots\ddots&\vdots\\ y_{O1}&\cdots\cdots\cdots&y_{OH}\end{bmatrix}\quad\begin{bmatrix}O:\text{Number of variables}\\ H:\text{Number of posterior samples}\end{bmatrix}$
Appendix C Additional Results
Here, we present the complete results analogous to Tables 1 and 2 for the larger test cases: ‘case500’ and ‘case2000’ with 500 and 2000 nodes, respectively, in Tables 6 and 7. In general, it is clear from the tables that the approaches presented in this paper outperform the state-of-the-art DNN-based approaches when limited training time and compute resources are provided. Between the supervised and the sandwich BNN models, it appears that there is no clear winner. Table 6 indicates that the supervised BNN outperforms the sandwich BNN model on the Max. Ineq. gap metric for the ‘case500’ whereas the trend is reversed for the case2000’ in Table 7. This minor variation is attributed to the fact that the maximum training time of $T_{\max}=600$ sec. is insufficient for both cases, and with more training time, these trends should look similar to the ones in Tables 1 and 2.
Another critical observation for both cases is that the ‘Max. Eq.’ value is substantially high even for the proposed best model (in Tables 6 and 7, respectively). However, it is clear that the proposed BNN proxies are better than standard DNN models with an order-of-magnitude difference. On the surface, readers may dismiss the efficacy of the proposed BNN models due to these large values. Still, these values have to be examined in conjunction with the error bounds on the voltage magnitude and voltage angles in Fig. 6. Examined together, the BNN model predictions have very low errors for the voltage magnitude and angles, and it is very much possible to develop a computationally inexpensive projection algorithm that projects this prediction onto the feasibility set of the ACOPF problem on lines of Zamzam & Baker (2020). This procedure would further reduce the Max. Eq. values and makes the projected solution usable. This is a minor detail regularly tackled in a power grid context and is not dealt with extensively in this paper.
@@ Appendix C Additional Results
Figure 6: Comparison of voltage magnitude and voltage angle error bounds (in logarithmic scale) across bus indices for case500. The plot illustrates that PCBs using theoretical Bernstein bounds with 2 $\times$ MPV from hypothesis equation 5 are tightest among all PCBs. We consider $\delta=0.95$ and 1000 out-of-sample testing data points i.e. $M=1000$.
Appendix D Concentration Bounds

### v3

@@ 4 Probabilistic Confidence Bounds
5 Numerical Results and Discussion
We test the proposed method on the Alternating Current Optimal Power Flow (ACOPF) problem, essential for the economic operation of electrical power grids (Molzahn et al., 2019). Efficient ACOPF proxies can mitigate climate change by enabling higher renewable energy integration, improving system efficiency by minimizing losses and emissions, and enhancing grid resiliency against extreme weather conditions (Rolnick et al., 2022a). ACOPF is a constrained optimization problem with nonlinear equality constraints and double-sided inequality bounds. It aims to find the most cost-effective generator set points while satisfying demand and adhering to physical and engineering constraints. The inputs are active and reactive power demands; outputs include generator settings, voltage magnitudes, and phase angles at each bus. We adopt the standard ACOPF formulation (Babaeinejadsarookolaee et al., 2019; Park & Van Hentenryck, 2023; Coffrin et al., 2018) and benchmark our method using the open-source OPFDataset from Torch_Geometric, which provides numerous solved ACOPF instances (see (Lovett et al., 2024)).
In our results, ‘Gap%’ denotes the average relative optimality gap compared to the objective values in the labeled testing instances. ‘Max Eq.’ and ‘Mean Eq.’ represent the maximum and mean equality gaps over all equality constraints, while ‘Max Ineq.’ and ‘Mean Ineq.’ indicate the same for inequality gaps in per unit, all averaged over testing instances. We compare our proposed method with the following state-of-the-art baseline supervised learning models available in the literature using the same labeled dataset and training time constraints, utilizing AI4OPT’s ML4OPF package (AI4OPT, 2023) (network and hyper-parameter details are in Appendix B):
  • Naïve MAE and Naïve MSE (Supervised): Use $l_{1}$-norm and $l_{2}$-norm loss functions, respectively, to measure differences between predicted and actual optimal solutions (Park & Van Hentenryck, 2023), incorporating a bound repair layer with a sigmoid function. The bound repair layer ensures that inequality constraints are always satisfied.
  • MAE + Penalty, MSE + Penalty, and LD (Supervised): Add penalty terms for constraint violations to the naïve MAE or MSE loss functions (Park & Van Hentenryck, 2023). The Lagrangian Duality (LD) method applies the $l_{1}$-norm as outlined in (Fioretto et al., 2020; Park & Van Hentenryck, 2023), and also uses a bound repair layer with a sigmoid function.
We exclude self-supervised constrained optimization methods like Primal-Dual Learning (PDL) (Park & Van Hentenryck, 2023) and DC3 (Donti et al., 2021) due to their significantly higher training times and computational demands, which violate the premise of this paper. For example, PDL requires over 125 minutes of training time for the ACOPF problem on a 118-node power network using a Tesla RTX6000 GPU. Methods that require solving alternating current power flow to recover solutions (e.g., (Zamzam & Baker, 2020)) are also excluded, as they result in high prediction times compared to DNN or BNN forward passes555see prediction time studies in (Donti et al., 2021) which suggest 10 times higher prediction time with power flow based projections (0̃.080 sec. compared to 0.001 sec. for DNN and 0.003 sec. per testing instance for proposed BNNs).. Furthermore, Graph Neural Network-based large models, such as (Piloto et al., 2024), require extensive training datasets—for instance, (Piloto et al., 2024) utilizes 270k training samples.
@@ 5 Numerical Results and Discussion
Tables 1 and 2 present the comparative performance of different methods for solving the ACOPF problem on the ‘case57’ and ‘case118’ test cases, containing 57 and 118 nodes, respectively. For ’case57’, the Sandwich BNN SvP method achieves the best Max Eq. performance (0.027), outperforming all other methods, including the standard Sandwich BNN, without compromising other metrics. All the methods proposed in this paper outperform the best DNN results, typically achieved with the LD+MAE model (last row in the tables). Similar trends are observed for the ‘case118’ as well. [^6] See Appendix C for similar tabular results and discussion of trends on larger test instances. It is important to contextualize the significance of these numerical improvements. In ACOPF problems, cost values are in USD, with the mean cost for ‘case118’ being \$97,000 or 9.7 in the per-unit system. Therefore, a 1% Gap corresponds to an expected difference of \$970 across the testing instances. A ‘Max Eq.’ value of 0.08 implies a maximum expected power imbalance of 8.0 Megawatts among all 118 nodes of ‘case118’. Thus, reducing the Max Eq. from 1.284 with the LD+MAE model to 0.089 with our Sandwich BNN SvP model represents a significant improvement.
Sandwich DNN in these tables is a DNN with the same network architecture as the BNN, trained under the same time constraints. This network is trained by alternating between minimizing a supervised MSE loss and and an unsupervised feasibility loss as in (2). In both these rounds, an $\ell_{1}$ weight regularization is added to prevent over-fitting. We see that this DNN model shows performance comparable to other DNN approaches. This is a strong indication that the superior performance of the sandwich BNN model is largely due to the superior performance of the BNN approach in the low data regime.
@@
Figure 4: Comparison of voltage magnitude and voltage angle error bounds (in logarithmic scale) across bus indices for ‘case57’ (top row) and ‘case118’ (bottom row). The plot illustrates that PCBs using theoretical Bernstein bounds with 2 $\times$ MPV from hypothesis (5) are tightest among all PCBs. We consider $\delta=0.95$ and 1000 out-of-sample testing data points i.e. $M=1000$.
@@
Figure 2 illustrates the growth of various metrics with increasing system size while keeping training resources constant. The proposed BNN methods exhibit significantly lower scaling in the expected maximum equality gap. Although Sandwich BNN SvP shows slightly higher scaling inequality gaps, the relative improvement in Max Eq. far outweighs these minor drawbacks. Notably, for ‘case500’ (see Table 8 in the Appendix C), the expected maximum power imbalance is below 0.5% of the mean real power demand ($1.7\times 10^{4}$ MW) and power grids already are equipped with spinning reserves (see (Ela et al., 2011)) that have reserve capacity to handle these imbalances. This is a significant improvement over the DNN models which have much higher ‘Mean Eq.’ values. Moreover, the ‘Max Ineq.’ growth could be easily suppressed by incorporating bound repair layers, as used in DNN models in ML4OPF (AI4OPT, 2023).
We perform robustness study on ‘case118’, by running five different trials of 10-min and 15-min training time with 512,1024, and 5048 supervised training samples keeping 2048 unsupervised samples constant across models and trials (See Appendix C for detailed graphs and tables). The results clearly shows that increasing training time and number of supervised samples decreases decreases %Gap, Max Eq. and Max Ineq., consistently, and keep Mean Eq. one order of magnitude lower and Mean Ineq. at order $10^{-5}$. The Sandwich BNN SvP model, provides 1.50% Gap, 0.094 as Max Eq. and 0.014 as Max Ineq. with 512 supervised training samples and 10-Min training time,, on average over five trials. The same model provides 1.46% Gap, 0.080 Max Eq. and 0.011 as Max Ineq. with 2048 supervised training samples and 15-Min Training time. This implies that five minute increase in training time with four times more training samples worth 2.66% reduction in optimality error (%Gap), 17.5% reduction in equality constraint feasibility error (Max Eq.) and more than 27% reduction in inequality constraint feasibility error (Max Ineq.).
This detailed study also highlight that smaller training times are not good enough to extract complete information from larger training datasets as various models with 2048 supervised samples perform better with 15-min training come compared to 10-min training time. For instance, Max Eq. is 6.18% lesser for 2048 samples when we increase training time from 10-Min to 15-Min, and %Gap values reduces from 1.52 to 1.46, on average. Lastly, we also observe that variability in training quality reduces with increasing number of supervised samples and training time.
@@
Next, we present results for Proabilistic Confidence bounds, described in Section 4. Figure 3 shows that $2\cdot\text{MPV}$ consistently serves as an empirical upper bound for total variance in error, validating its robustness across models and system configurations [^7] Total variance in error is assumed to stabilizing with 1000 testing samples (see Figure 7 in Appendix C).. In Figure 4, the theoretical Bernstein bounds using $\mathbb{V}_{e}=2\cdot\text{MPV}$ provide tight, practical bounds, whereas Hoeffding’s bounds are overly conservative and not useful for grid operations. For example, in ‘case118’, the Bernstein bound ensures a probabilistic guarantee on voltage constraint satisfaction, such that a predicted voltage between 0.91–1.09 pu guarantees no violations within the ACOPF limits of 0.90–1.10 pu, i.e., the maximum value of the Bernstein bound on the error is 0.010 pu across all nodes. Compared to Hoeffding’s bound ($0.064$ pu) or the empirical Bernstein bound ($0.018$ pu), the Bernstein bound ($0.010$ pu) is far tighter and more practical, highlighting the benefits of BNNs for optimization proxies. The error bounds for the ‘case500’ is provided in Fig. 7 of the Appendix C. Finally, we note that both Hoeffding’s and empirical Bernstein bounds can also be obtained by testing DNN models across $M$ samples [^8] A form of MPV can be obtained via Ensemble DNNs (Ganaie et al., 2022), however, it will lead to very high computational requirement compared to the BNN..
@@ 6 Conclusions
Reproducibility
We use the open-source ACOPF datasets, provided with Torch_geometric (Lovett et al., 2024), to train and test our models as well as standard DNN models. Furthermore, the beginning of Section 5 provides details of the DNN models used to compare the performance of the proposed methods. These models are available in AI4OPT’s open-source ML4OPF package (AI4OPT, 2023). Hyper-parameter details for these models and the proposed methods are provided in Appendix B. The code used in this paper can be found at https://github.com/kaarthiksundar/BNN-OPF/. Additional experimental results can be found in the Supplementary Information.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. In particular, enhancing the solution of optimization problems will result in more efficient resource utilization, helping industries lower costs and reduce environmental impact. Additionally, refining the ACOPF solution pipeline will play a crucial role in addressing climate change by optimizing renewable energy usage and ensuring the reliable operation of the power grid (Rolnick et al., 2022b).
@@ Appendix A ACOPF Problem: Modeling and Dataset
The alternating current optimal power flow (ACOPF) problem is essential for power grid operations and planning across various time scales. It determines generator set-points for real and reactive power that minimize generation costs while meeting power demand and satisfying physical and operational constraints on output variables. We follow the ACOPF model given in PowerModels (Coffrin et al., 2018), and take the dataset from Torch_geometric ((Lovett et al., 2024)).
@@ B.2.1 ML4OPF:
All other hyperparameters are set at default ML4OPF values (AI4OPT, 2023).
@@ B.3 Structure of posterior prediction matrix (PPM)
$\Large\mathbf{Y}\equiv\begin{bmatrix}y_{11}&\cdots\cdots\cdots&y_{1H}\\ \vdots&\ddots\ddots\ddots&\vdots\\ \cdots&y_{\texttt{Variable,Sample}}&\cdots\\ \vdots&\ddots\ddots\ddots&\vdots\\ y_{O1}&\cdots\cdots\cdots&y_{OH}\end{bmatrix}\quad\normalsize\begin{bmatrix}O:\text{Number of variables}\\ H:\text{Number of posterior samples}\end{bmatrix}$
Appendix C Robustness Study on 118-System
Figure 5: Summary of Results with more variations in supervised data, for higher training (15 min) for sandwich and supervised models with SvP. Note that this computation was done on a Max Mini M4 with 24GB RAM.
Table 6: Comparison of Average Values (Over Five Trials) for BNN Models: 10/15 min, Sandwich SvP/Supervised SvP. Note that this computation was done on a Max Mini M4 with 24GB RAM.
N Sandwich BNN SvP, 15-Min Supervised BNN SvP, 15-Min
%Cost Max Eq Mean Eq Max Ineq Mean Ineq %Cost Max Eq Mean Eq Max Ineq Mean Ineq
512 1.5305 0.1041 0.01998 0.01285 0.000046 1.5904 0.1755 0.02403 0.01446 0.000060
1024 1.5097 0.08948 0.01872 0.01442 0.000049 1.5615 0.1201 0.02075 0.01892 0.000075
2048 1.4691 0.08005 0.01709 0.01170 0.000042 1.5534 0.08776 0.01839 0.01906 0.000073
N Sandwich BNN SvP, 10-Min Supervised BNN SvP, 10-Min
%Cost Max Eq Mean Eq Max Ineq Mean Ineq %Cost Max Eq Mean Eq Max Ineq Mean Ineq
512 1.5071 0.09438 0.01912 0.01453 0.000054 1.5807 0.1644 0.02323 0.01524 0.000061
1024 1.5194 0.08693 0.01865 0.01415 0.000047 1.5562 0.1108 0.02025 0.01679 0.000065
2048 1.5297 0.08532 0.01855 0.01288 0.000042 1.5387 0.08686 0.01843 0.01912 0.000072
Table 7: Comparison of Average Values (Over Five Trials) for BNN models: 10/15-min, Sandwich/Supervised. Note that this computation was done on a Max Mini M4 with 24GB RAM.
N Sandwich BNN, 15 Min Supervised BNN, 15 Min
%Cost Max Eq Mean Eq Max Ineq Mean Ineq %Cost Max Eq Mean Eq Max Ineq Mean Ineq
512 1.5298 0.1349 0.01810 0.01255 0.000045 1.5905 0.2281 0.02318 0.01425 0.000060
1024 1.5074 0.1106 0.01661 0.01407 0.000048 1.5594 0.1609 0.01957 0.01868 0.000074
2048 1.4687 0.09156 0.01474 0.01135 0.000041 1.5540 0.1114 0.01690 0.01890 0.000072
N Sandwich BNN, 10-Min Supervised BNN, 10-Min
%Cost Max Eq Mean Eq Max Ineq Mean Ineq %Cost Max Eq Mean Eq Max Ineq Mean Ineq
512 1.5054 0.1194 0.01710 0.01426 0.000053 1.5780 0.2204 0.02215 0.01502 0.000060
1024 1.5188 0.1064 0.01654 0.01376 0.000046 1.5573 0.1499 0.01888 0.01652 0.000064
2048 1.5284 0.09725 0.01623 0.01251 0.000041 1.5377 0.1081 0.01664 0.01887 0.000071
Figure 5 shows results of 5-Trials of proposed method of learning for three key metrics—Gap, Max Eq., and Max Ineq.—across different sample sizes (N = 512, 1024, 2048) for ‘case118’ on a Mac Mini M4 with 24GB RAM. The red dashed lines with circular markers represent the mean values obtained from the sandwich training approach, while the blue solid lines with square markers correspond to the mean values from the supervised training approach. Error bars depict the minimum and maximum observed values for each method, providing insight into the variability of results. From the figure, we observe that as the number of samples increases, both methods exhibit a trend of increasing Max Eq. and Max Ineq. values, while % Gap remains relatively stable.
The supervised approach generally achieves lower variability (shorter error bars), indicating a more consistent performance. The sandwich method, while slightly less stable, maintains comparable mean values across the three metrics. These results suggest that supervised learning provides more predictable outcomes, whereas the sandwich method may introduce greater flexibility with minor variations in performance. Detailed results and and graphs are given in Table 7, Table 6, Figure 11 to Figure 15.
Results in Figure 11 to 15 are of five trials with identical hyperparameters demonstrate consistent improvement across all parameters except for max inequality. Max inequality remains challenging to suppress further, as it is already one order of magnitude better than max equality. Also the spread of error decreases with more samples indicating robust learning performance with more data. Note that these computations was done on a Mac Mini M4 with 24GB RAM.
C.1 Larger System Results
@@
Here, we present the complete results analogous to Tables 1 and 2 for the larger test cases: ‘case500’ and ‘case2000’ with 500 and 2000 nodes, respectively, in Tables 8 and 9. In general, it is clear from the tables that the approaches presented in this paper outperform the state-of-the-art DNN-based approaches when limited training time and compute resources are provided. Between the supervised and the sandwich BNN models, it appears that there is no clear winner. Table 8 indicates that the supervised BNN outperforms the sandwich BNN model on the Max. Ineq. gap metric for the ‘case500’ whereas the trend is reversed for the case2000’ in Table 9. This minor variation is attributed to the fact that the maximum training time of $T_{\max}=600$ sec. is insufficient for both cases, and with more training time, these trends should look similar to the ones in Tables 1 and 2.
Another critical observation for both cases is that the ‘Max. Eq.’ value is substantially high even for the proposed best model (in Tables 8 and 9, respectively). However, it is clear that the proposed BNN proxies are better than standard DNN models with an order-of-magnitude difference. On the surface, readers may dismiss the efficacy of the proposed BNN models due to these large values. Still, these values have to be examined in conjunction with the error bounds on the voltage magnitude and voltage angles in Fig. 7. Examined together, the BNN model predictions have very low errors for the voltage magnitude and angles, and it is very much possible to develop a computationally inexpensive projection algorithm that projects this prediction onto the feasibility set of the ACOPF problem on lines of (Zamzam & Baker, 2020). This procedure would further reduce the Max. Eq. values and makes the projected solution usable. This is a minor detail regularly tackled in a power grid context and is not dealt with extensively in this paper.
@@
Figure 7: Comparison of voltage magnitude and voltage angle error bounds (in logarithmic scale) across bus indices for case500. The plot illustrates that PCBs using theoretical Bernstein bounds with 2 $\times$ MPV from hypothesis (5) are tightest among all PCBs. We consider $\delta=0.95$ and 1000 out-of-sample testing data points i.e. $M=1000$.
Appendix D Performance on a non-OPF dataset
The semi-supervised training approach described in this work can be used to train optimization proxies for any constrained optimization problem over continuous variables. In this section we look at the performance of our method on the following non-convex optimization problem described in Donti et.al. (Donti et al., 2021).
$\min_{y\in\mathbb{R}^{n}}\;\frac{1}{2}y^{T}Qy+p^{T}\sin(y),\quad\text{s.t. }Ay=x,\;Gy\leq h.$ (6)
The sinusoid function in the objective makes this a non-convex problem. We used the code provided by Donti et.al. to generate two sets of randomized instances of this problem. The comparative performance of the sandwiched training method against purely supervised BNN training is given in Table 10
Model (nV, nEq, nInEq) Max-Eq violation Max-InEq violation GapTime
BNN-Supervised (70,20,50) 0.484 0.170 2.6 800s
BNN-Sandwich (70,20,50) 0.298 0.109 7.1 800s
BNN-Supervised (20,10,20) 0.264 0.108 1.4 400s
BNN-Sandwich (20,10,20) 0.114 0.000 0.07 400s
Table 10: nV, nEq and nInEq are respectively the number of variables, equality constraints and inequality constraints in the optimization problem. For these experiments we use $2^{12}$ samples in the supervised stage and $2^{9}$ samples in the unsupervised stage. Results above are evaluated on 100 testing instances and averaged over three different random instantiations of the training procedure. These show the advantages of using the sandwiched approach for training a BNN model in the low-data/time-constrained regime for this problem. . The hardware used here is the Google Collab instance with Intel(R) Xeon(R) CPU @ 2.20GHz.
Appendix E Concentration Bounds